Find the first four terms and the 100th term of the sequence whose nth term is given. an = 4n - 2 a1 az a 3 a 100
Answer #1

Given that an=40-2 puf n=1 ) a = 40-2. - 460-254-23 2 n=2 a = 4(2) 228-2=6 n=3 a3 = 4(3) 2 = /2-2 =10 na47 a4=4(4)-2 = 16-2=1